OpenStudy (anonymous):

What is 1/2 times 2/3

OpenStudy (anonymous):

1/2 * 2/3 = (1*2) / (2*3)=2/6=1/3

OpenStudy (anonymous):

Shortcut: 1 2 1 - * - (the 2s cancel out, and it equals,) -- 2 3 3

OpenStudy (anonymous):

Thanks I couldn't remember how to work the problem.

OpenStudy (anonymous):

Yw